OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (radar):

rewrite the line in slope intercept form y=(4/3)x+4 so the slope is a 4/3 and so will the slope of the line that will be parallel.

OpenStudy (anonymous):

thanks

OpenStudy (radar):

The given point is (-3,4) so we use the slope formula again: 4 y-4 --- = ----- 3 x-(-3) 4x+12=3y-12 3y=4x + 24 y=(4/3)(x)+8

OpenStudy (radar):

that is the equation for the parallel line, parallel to4x-3y=-12 and passes through (-3,4)